The length of a bicycle pedal arm is 0.142m, and a downward


Force of 111N is applied to the pedal by the rider's foot. What is the magnitude of the torque about the pedal arm pivot when the arm makes an angle of


(A.) 30°


(B) 90°


(C.) 180°


With the vertical?

1
Expert's answer
2021-12-10T13:20:59-0500

(A) "\tau=F\cdot l\cdot\sin30°=111\cdot0.142\cdot\sin30°=7.9\ (N\cdot m)"


(B) "\tau=F\cdot l\cdot\sin90°=111\cdot0.142\cdot\sin90°=15.8\ (N\cdot m)"


(C) "\tau=F\cdot l\cdot\sin180°=111\cdot0.142\cdot\sin180°=0\ (N\cdot m)"